# Semi-empirical constraints on the HI mass function of star-forming galaxies and $\Omega_{\rm HI}$ at $z\sim 0.37$ from interferometric surveys

### v1

@@ 4.1 The reference stellar mass function
As anticipated in Sect. 2.1, we adopt the stellar mass function for star-forming galaxies at $0.50.2<z<$ 0.5 from the COSMOS2020 photometric catalog (Weaver et al., 2023). In particular, we take advantage of the publicly-available posterior distributions of the parameters for the fitted double Schechter profile as follows.
@@ 4.2 Converting stellar masses into HI masses
To pass from the stellar mass sample obtained as described in Section 4.1, we adopt a scaling relation from literature results. In particular, we adopt the most recent findings from Bianchetti et al. (2025). As described in Sect. 2.3, therein the authors performed H I stacking in $4$ stellar mass bins for the combined case, and $3$ stellar mass separately for MIGHTEE and CHILES. For the different cases, the authors fitted a single power law model of the type $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=b\log_{10}(M_{\star}/{\rm M}_{\odot})+c$, with $b$ and $c$ free parameters, i.e. a linear function in logarithmic space (hereafter the linear model). In this work, we generalize this framework by fitting also a second-order polynomial in logarithmic space, parametrized as $\log_{10}(M_{\rm HI})=a\log_{10}(M_{\star})^{2}+b\log_{10}(M_{\star})+c$ (hereafter the polynomial model). This higher-order polynomial allows to capture a potential deviation from a single power law model — i.e. a linear model in logarithmic space — which has been shown to be the case for results at $z\sim 0$ based on direct H I detection (see e.g. Huang et al., 2012; Maddox et al., 2015; Parkash et al., 2018). Such a phenomenon is typically described by means of a functional form which decreases more quickly towards low $M_{\rm HI}$ and flattens toward high $M_{\rm HI}$ values. To account for this behaviour, the following double power law model was proposed (e.g. Maddox et al., 2015; Parkash et al., 2018; Pan et al., 2023):
@@
$\log_{10}(M_{\rm HI})=\log_{10}\left(\frac{M_{0}}{\left(\frac{M_{\star}}{M_{\rm tr}}\right)^{\alpha}+\left(\frac{M_{\star}}{M_{\rm tr}}\right)^{\beta}}\right)\quad,$ (4)
@@
with $M_{0}$, $M_{\rm tr}$, $\alpha$ and $\beta$ free parameters. We notice that $\alpha$ and $\beta$ denote the low- and high-mass end slopes and that this model reduces to the single power law model when $\alpha=\beta$. In this work, since we have at most $4$ data points available for the combined stacking case, we cannot fit this model — which has $4$ free parameters — but at most a model with $3$ free parameters, otherwise we would be overfitting our data. This motivates further the choice of a second-order polynomial, as alternative to a linear fit. Later on in the paper, we will show that the best-fitting polynomial model qualitatively reproduces the same trend as the double power law model discussed above. We notice that for the cases of the stacking performed separately on MIGHTEE and CHILES, for which we have only $3$ data points each, we are again in a situation where a full fit of a second-order polynomial with $3$ free parameter would result into overfitting. We will discuss this issue in Section 5.1 and show how we reduce the number of free parameters from $3$ to $2$ to have a well-defined fit also in the MIGHTEE and CHILES separate cases.
@@
We estimate the best-fitting parameters for the linear and polynomial models and the related uncertainties by means of a bootstrap resampling (see Bianchetti et al. (2025)). While this has already been performed by Bianchetti et al. (2025) for the linear model, it had not been done for the polynomial one. In particular, we resample the data points underlying the fit by randomly-sampling a lognormal uncertainty based on the uncertainty on the estimation of $M_{\rm HI}$ of each data point. [^3] We caution that the uncertainty on the $M_{\rm HI}$ estimates, obtained via jackknife resampling by Bianchetti et al. (2025), refer to the error on the measurement and they do not represent the full underlying $M_{\rm HI}$ distribution. In particular, the jackknife uncertainty typically underestimates the true dispersion. For each bootstrap sample, we then apply the full forward model and obtain the HIMF.
@@
In this work, since we are using a different scaling relation, we also need to assume a different value for the scatter. In particular, we make the same assumption as the one from Bera et al. (2022) and use the value of the scatter at $z\sim 0$ from the $M_{\rm HI}-M_{\star}$ scaling relation measured by the xGASS survey (Catinella et al., 2018), i.e. $\sigma=0.39$ dex. We study the specific impact of the scatter in our framework in Appendix B, where we show that the value for the scatter largely affects the high-mass end of the HIMF. Specifically, the high-mass end of the HIMF is shown to increase with increasing value of the scatter. Therefore, we incorporate an additional source of uncertainty on the scatter as follows. For each step of our Monte Carlo procedure, we add to the scatter an error randomly-sampled from a Gaussian distribution with zero mean and standard deviation $\sigma_{\rm err}=0.05$, conservatively covering a much wider range of credible values that the scatter can assume. In this way, we self-consistently account for the error on the scatter in each of the sampling steps and propagate the related uncertainty throughout our forward model. In Appendix B, we show that this source of stochasticity yields no significant difference on the uncertainties on the best-fitting parameters of the HIMF with respect to the case where no scatter is assumed, nor introduces any obvious systematic shift in the median values. Therefore, in the remainder of the paper we assume as fiducial configuration the one including the error on the scatter, even though we argue that this has been shown to have little impact.
@@ 4.3 Fitting the HI mass function and derivation of $\Omega_{\rm HI}$
In order to perform an unbiased estimate of the best-fitting parameters, we first establish the minimum $M_{\rm HI}$ that we can trust given our model, which is equivalent to the observed completeness limit. While it is not trivial to establish such a limit as classical methods would not be applicable in the present case, we limit ourselves to a visual inspection of the resulting HIMF. We discuss this in Appendix D. Fig. 8 reveals that at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\lesssim 9$ the HIMF shows clear signs of incompleteness. Based on this visual estimate, we conservatively fit the Schechter model only for $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.2$.
@@
$\Phi(M_{\rm HI})=\log(10)\leavevmode\nobreak\ \tilde{\Phi}\left(\frac{M_{\rm HI}}{\tilde{M}}\right)^{\alpha+1}e^{\left(-M_{\rm HI}/\tilde{M}\right)}$ (5)
where the free parameters $\alpha$, $\tilde{\Phi}$, and $\tilde{M}$ represent the faint-end slope, the normalization, and the ‘knee’ mass of the mass function. For the fitting procedure, we adopt a Markov Chain Monte Carlo approach and use the affine-invariant emcee ensemble sampler (Foreman-Mackey et al., 2013). For each of the tested models, we run $50,000$ steps and discard the first $1,000$ to conservatively avoid the burn-in phase. We then estimate the best-fitting value as the median ($50th$ percentile) and the associated uncertainty as difference ($84th-50th$) percentile (upper) and ($50th-16th$) percentile (lower).
@@ 5.1 Fitting the $M_{\rm HI}-M_{\star}$ relation
From Fig. 1, we see that both the linear and the polynomial models provide reasonable fits to the data. In order to quantify which model best fits the data and what implications this has, we compute the ’reduced chi square’ $\chi^{2}/({\rm degrees\leavevmode\nobreak\ of\leavevmode\nobreak\ freedom})$ and report the values in the last column of Table 2. We notice that:
  • for all the three cases, the $\chi^{2}/{\rm d.o.f.}$ is much smaller for the polynomial fit than for the linear fit;
  • the parameters $a$ in the polynomial fit is found to be negative $a<0$ at a significance $>10\sigma$ in all the cases.
The first fact provides the first evidence for the preference of the polynomial model over the linear model beyond the Local universe. The second implies that, as anticipated in Section 4.2, a second-order polynomial with $a<0$ has the same qualitative properties as a double power law model, i.e. it decreases more rapidly than a linear model towards low H I masses and flattens towards high H I masses.
@@
MIGHTEE - $0.29\pm 0.03$ $7.03\pm 0.29$ $162.35$
CHILES - $0.33\pm 0.06$ $6.56\pm 0.60$ $25.68$
Combined - $0.33\pm 0.03$ $6.58\pm 0.33$ $76.29$
@@
MIGHTEE $-0.068\pm 0.002$ $1.65\pm 0.03$ $0.49$
CHILES $-0.065\pm 0.006$ $1.61\pm 0.06$ $1.59$
Combined $-0.063\pm 0.003$ $1.60\pm 0.04$ $1.16$
@@
Survey z $\log_{10}(\tilde{M})$ $\alpha$ $\log_{10}(\tilde{\Phi})$ $\Omega_{\rm HI}\times 10^{-4}$ Reference
ALFALFA.100 0 $9.93\pm 0.01\pm 0.005$ $-1.25\pm 0.02\pm 0.1$ $-2.28\pm 0.02\pm 0.07$ $3.90\pm 0.10$ Jones et al. (2018)
MIGHTEE $1/V_{\rm max}^{5S_{\rm lim}}$ 0.04 $10.04^{+0.24}_{-0.24}$ $-1.29^{+0.37}_{-0.26}$ $-2.29^{+0.32}_{-0.36}$ $5.26^{+0.91}_{-0.95}$ Ponomareva et al. (2023)
MIGHTEE $1/V_{\rm max}^{8S_{\rm lim}}$ 0.04 $10.07^{+0.22}_{-0.24}$ $-1.40^{+0.42}_{-0.24}$ $-2.31^{+0.35}_{-0.38}$ Ponomareva et al. (2023)
MIGHTEE MML 0.04 $10.19^{+0.10}_{-0.13}$ $-1.44^{+0.13}_{-0.10}$ $-2.47^{+0.19}_{-0.14}$ $6.08^{+0.30}_{-0.30}$ Ponomareva et al. (2023)
AUDS 0.16 $10.14\pm 0.09$ $-1.37\pm 0.05$ $-2.55\pm 0.01$ $3.55\pm 0.29$ Hoppmann et al. (2015)
BUDHIES 0.2 $9.77^{+0.16}_{-0.04}$ $-1.49\pm 0.48$ $-2.25\pm 0.03$ $4.10\pm 4.43$ Gogate (2022)
HI int. mapping 0.32 $9.46^{+0.38}_{-0.31}$ $-1.29^{+0.36}_{-0.39}$ $-1.80^{+0.65}_{-0.72}$ $4.13^{+5.55}_{-1.58}$ Paul et al. (2023)
Combined (poly.) 0.37 $10.16^{+0.08}_{-0.08}$ $-1.48^{+0.10}_{-0.09}$ $-2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$ This work
Combined (linear) 0.37 $10.07^{+0.08}_{-0.08}$ $-1.42^{+0.12}_{-0.11}$ $-2.19^{+0.13}_{-0.14}$ $8.51^{+0.69}_{-0.56}$ This work
Table 3: Numerical results for the best-fitting parameters to the HIMF using a Schechter model as in Eq. 5, for the results from this work and from different literature results. The first column reports the surveys the results refer to; the second column reports the median redshift at which the HIMF has been measured; the third column reports the sample size (if applicable), the fourth, fifth and sixth column report the best-fitting values for the $\log_{10}(\tilde{M})$, $\alpha$, and $\log_{10}(\tilde{\Phi})$; the seventh column reports the inferred value for $\Omega_{\rm HI}$; the eight column displays the reference presenting the results.
@@
Fig. 2 shows the resulting HIMF from this work as yellow stars (linear model) and blue squares (polynomial model). We display as blue and yellow solid lines the best-fitting Schechter profiles as found from the median of the posterior distributions of the parameters. We also display the results for the HIMF from the literature reported in Table 3. In particular, we show the results from ALFALFA at $z\sim 0$ (Jones et al., 2018, light green dashed), from MIGHTEE at $0.0840<z<$ 0.084 (Ponomareva et al., 2023, red dotted and gray dotted-dashed), from the AUDS (Hoppmann et al., 2015, brown dotted-dashed) and BUDHIES (Gogate, 2022, cyan dashed) surveys at $z\sim 0.16$ and $z\sim 0.2$, respectively (all based on direct detections), from a HI intensity mapping experiment at $z\sim 0.32$ (Paul et al., 2023, purple dotted-dashed), and green circles and orange the semi-empirical HIMF results obtained by Bera et al. (2022) at $z\sim 0.35$ and Chowdhury et al. (2024) at $z\sim 1$ following similar approaches to the one adopted herein. We find evolution between the HIMF reported in this work and the one at $z\sim 0$ from ALFALFA. In particular, the HIMF derived here with both models have a consistently higher normalization than the ALFALFA HIMF, on average by a factor $\sim 2$. Compared to the MIGHTEE results at lower redshift and to the AUDS survey at $z\sim 0.16$, the HIMF from this work has a larger normalization at low $M_{\rm HI}$ values, whereas they tend to converge towards high $M_{\rm HI}$. The results from BUDHIES, from Bera et al. (2022), and from Paul et al. (2023) have comparable normalization to the one of our results at low $M_{\rm HI}$, but they present a rapid fall-off towards high $M_{\rm HI}$, dropping below the ALFALFA HIMF at $\log_{10}(M_{\star}/{\rm M_{\odot}})\lesssim 9.6$. Essentially, there is evidence for evolution between our results at $z\sim 0.37$ and the results by Chowdhury et al. (2024) at $z\sim 1$, the latter having a consistently larger normalization at the probed H I masses.
@@
  • the resulting values for $\Omega_{\rm HI}$ obtained by assuming the linear and the polynomial model for the scaling relation differ at $\sim 1.1\sigma$ for the combined stacking;
@@
Figure 2: Compilation of HIMF results from this work and from the literature. We show as yellow stars and blue squares the resulting HIMF from the linear and polynomial models, respectively, as well as their best-fitting models as solid lines of the same color. We display the HIMF from ALFALFA at $z\sim 0$ (Jones et al., 2018) as a green dashed line, from MIGHTEE at $0.0840<z<$ 0.084 as red dotted ($1/V_{\rm max}$) and gray dotted-dashed (MML) lines (Ponomareva et al., 2023), from AUDS at $z\sim 0.16$ as a brown dotted-dashed line (Hoppmann et al., 2015), from BUDHIES at $z\sim 0.2$ as a cyan dashed line (Gogate, 2022), from the intensity mapping experiment by at $z\sim 0.32$ as a purple dotted-dashed line (Paul et al., 2023), and as green circles and orange diamonds the results from a similar approach as the one followed herein applied to stacking results from uGMRT data at $z\sim 0.35$ (Bera et al., 2022) and $z\sim 1$ (Chowdhury et al., 2024).

### v2

@@ 4.1 The reference stellar mass function
As anticipated in Sect. 2.1, we adopt the stellar mass function for star-forming galaxies at $0.2<z<0.5$ from the COSMOS2020 photometric catalog (Weaver et al., 2023). In particular, we take advantage of the publicly-available posterior distributions of the parameters for the fitted double Schechter profile as follows.
@@ 4.2 Converting stellar masses into HI masses
To pass from the stellar mass sample obtained as described in Section 4.1, we adopt a scaling relation from literature results. In particular, we adopt the most recent findings from Bianchetti et al. (2025). As described in Sect. 2.3, therein the authors performed H I stacking in $4$ stellar mass bins for the combined case, and $3$ stellar mass bins separately for MIGHTEE and CHILES. For the different cases, the authors fitted a single power law model of the type $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=b\log_{10}(M_{\star}/{\rm M}_{\odot})+c$, with $b$ and $c$ free parameters, i.e. a linear function in logarithmic space (hereafter the linear model). In this work, we generalize this framework by fitting also a second-order polynomial in logarithmic space, parametrized as $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=a\log_{10}(M_{\star}/{\rm M}_{\odot})^{2}+b\log_{10}(M_{\star}/{\rm M}_{\odot})+c$ (hereafter the polynomial model). This higher-order polynomial allows to capture a potential deviation from a single power law model — i.e. a linear model in logarithmic space — which has been shown to be the case for results at $z\sim 0$ based on direct H I detection (see e.g. Huang et al., 2012; Maddox et al., 2015; Parkash et al., 2018). Such a phenomenon is typically described by means of a functional form which decreases more quickly towards low $M_{\rm HI}$ and flattens toward high $M_{\rm HI}$ values. To account for this behaviour, the following double power law model was proposed (e.g. Maddox et al., 2015; Parkash et al., 2018; Pan et al., 2023):
@@
$\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=\log_{10}\left(\frac{M_{0}/{\rm M}_{\odot}}{\left(\frac{M_{\star}}{M_{\rm tr}}\right)^{\alpha}+\left(\frac{M_{\star}}{M_{\rm tr}}\right)^{\beta}}\right)\quad,$ (2)
@@
with $M_{0}$, $M_{\rm tr}$, $\alpha$ and $\beta$ free parameters. We notice that $\alpha$ and $\beta$ denote the low- and high-mass end slopes and that this model reduces to the single power law model when $\alpha=\beta$. In this work, since we have at most $4$ data points available for the combined stacking case, we cannot fit this model — which has $4$ free parameters — but at most a model with $3$ free parameters, otherwise we would be overfitting our data. This motivates further the choice of a second-order polynomial, as alternative to a linear fit. Later on in the paper, we will show that the best-fitting polynomial model qualitatively reproduces the same trend as the double power law model discussed above. We notice that for the cases of the stacking performed separately on MIGHTEE and CHILES, for which we have only $3$ data points each, we are again in a situation where a full fit of a second-order polynomial with $3$ free parameter would result in overfitting. We will discuss this issue in Section 5.1 and show how we reduce the number of free parameters from $3$ to $2$ to have a well-defined fit also in the MIGHTEE and CHILES separate cases.
@@
We estimate the best-fitting parameters for the linear and polynomial models and the related uncertainties by means of a bootstrap resampling (see Bianchetti et al. (2025)). While this has already been performed by Bianchetti et al. (2025) for the linear model, it had not been done for the polynomial one. In particular, we resample the data points underlying the fit by randomly-sampling a lognormal uncertainty based on the uncertainty on the estimation of $M_{\rm HI}$ of each data point. [^3] We caution that the uncertainties on the $M_{\rm HI}$ estimates, obtained via jackknife resampling by Bianchetti et al. (2025), refer to the error on the measurement and they do not represent the full underlying $M_{\rm HI}$ distribution. In particular, the jackknife uncertainty typically underestimates the true dispersion. For each bootstrap sample, we then apply the full forward model and obtain the HIMF.
@@
In this work, since we are using a different scaling relation, we also need to assume a different value for the scatter. In particular, we make the same assumption as the one from Bera et al. (2022) and use the value of the scatter at $z\sim 0$ from the $M_{\rm HI}-M_{\star}$ scaling relation measured by the xGASS survey (Catinella et al., 2018), i.e. $\sigma=0.39$ dex. We study the specific impact of the scatter in our framework in Appendix B, where we show that the value for the scatter largely affects the high-mass end of the HIMF. Specifically, the high-mass end of the HIMF is shown to increase with increasing value of the scatter. Therefore, we incorporate an additional source of uncertainty on the scatter as follows. For each step of our Monte Carlo procedure, we add to the scatter an error randomly-sampled from a Gaussian distribution with zero mean and standard deviation $\sigma_{\rm err}=0.05$. The value for this additional uncertainty contribution is motivated by covering the range of values $\sigma=(0.39\pm 0.05)$ dex, i.e. a conservatively wide range of credible values that the scatter can assume according to $z=0$ observations. In this way, we self-consistently account for the error on the scatter in each of the sampling steps and propagate the related uncertainty throughout our forward model. In Appendix B, we show that this source of stochasticity yields no significant difference on the uncertainties on the best-fitting parameters of the HIMF with respect to the case where no scatter is assumed, nor introduces any obvious systematic shift in the median values. Therefore, in the remainder of the paper we assume as fiducial configuration the one including the error on the scatter, even though we argue that this has been shown to have little impact. In addition, in Appendix C we study the impact of relaxing the assumption of a Gaussian symmetric scatter in log space, and test whether adopting a more realistic asymmetric functional form has any impact on the final results. Specifically, we model the scatter as a lognormal distribution in log space with parameters determined by fitting the true H I distribution at fixed $M_{\star}$ as measured by the cross-matched SDSS-ALFALFA catalog Durbala et al. (2020). We find that accounting for an asymmetric scatter has no significant impact with respect to assuming a symmetric scatter, and stick to the simpler Gaussian model for simplicity.
@@ 4.3 Fitting the HI mass function and derivation of $\Omega_{\rm HI}$
In order to perform an unbiased estimate of the best-fitting parameters, we first establish the minimum $M_{\rm HI}$ that we can trust given our model, which is equivalent to the observed completeness limit. While it is not trivial to establish such a limit as classical methods would not be applicable in the present case, we limit ourselves to a visual inspection of the resulting HIMF. We discuss this in Appendix D. Fig. 8 reveals that at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\lesssim 9$ the HIMF shows clear signs of incompleteness. Based on this visual estimate, we conservatively fit the Schechter model only at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.2$ for the polynomial model and at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.4$ for the linear model.
@@
$\Phi(M_{\rm HI})=\log(10)\penalty 10000\ \tilde{\Phi}\left(\frac{M_{\rm HI}}{\tilde{M}}\right)^{\alpha+1}e^{\left(-M_{\rm HI}/\tilde{M}\right)}$ (3)
where the free parameters $\alpha$, $\tilde{\Phi}$, and $\tilde{M}$ represent the faint-end slope, the normalization, and the ‘knee’ mass of the mass function. For the fitting procedure, we adopt a Markov Chain Monte Carlo approach and use the affine-invariant emcee ensemble sampler (Foreman-Mackey et al., 2013). For each of the tested models, we run $50,000$ steps and discard the first $1,000$ to conservatively avoid the burn-in phase. We then estimate the best-fitting value as the median ($50th$ percentile) and the associated uncertainty as difference (84th50th84th-50th) percentile (upper) and (50th16th50th-16th) percentile (lower).
@@ 5.1 Fitting the $M_{\rm HI}-M_{\star}$ relation
From Fig. 1, we see that both the linear and the polynomial models provide reasonable fits to the data. We compute the ’reduced chi square’, $\chi^{2}/({\rm degrees\penalty 10000\ of\penalty 10000\ freedom})$, and report the values in the last column of Table 2. Given the low number of degrees of freedom, it is hard to meaningfully establish which among the two models is preferred. Therefore, we decide to adopt as fiducial model the polynomial one which, as anticipated in Section 4.2, having $a<0$ has the same qualitative properties as a double power law model, i.e. it decreases more rapidly than a linear model towards low H I masses and flattens towards high H I masses.
@@
MIGHTEE - $0.286\pm 0.032$ $7.029\pm 0.290$ $2.764$
CHILES - $0.328\pm 0.065$ $6.547\pm 0.607$ $0.174$
Combined - $0.326\pm 0.030$ $6.579\pm 0.339$ $2.670$
@@
MIGHTEE $0.070\pm 0.002$ $1.693\pm 0.027$ - $0.005$
CHILES $0.067\pm 0.006$ $1.651\pm 0.061$ - $1.043$
Combined $0.066\pm 0.003$ $1.648\pm 0.064$ - $1.002$
@@
Survey z $\log_{10}(\tilde{M}/{\rm M}_{\odot})$ $\alpha$ $\log_{10}(\tilde{\Phi}/{\rm Mpc}^{-3})$ $\Omega_{\rm HI}\times 10^{-4}$ Reference
ALFALFA.100 0 $9.93\pm 0.01\pm 0.005$ $1.25\pm 0.02\pm 0.1$ $2.28\pm 0.02\pm 0.07$ $3.90\pm 0.10$ Jones et al. (2018)
MIGHTEE $1/V_{\rm max}^{5S_{\rm lim}}$ 0.04 $10.04^{+0.24}_{-0.24}$ $1.29^{+0.37}_{-0.26}$ $2.29^{+0.32}_{-0.36}$ $5.26^{+0.91}_{-0.95}$ Ponomareva et al. (2023)
MIGHTEE $1/V_{\rm max}^{8S_{\rm lim}}$ 0.04 $10.07^{+0.22}_{-0.24}$ $1.40^{+0.42}_{-0.24}$ $2.31^{+0.35}_{-0.38} -$ Ponomareva et al. (2023)
MIGHTEE MML 0.04 $10.19^{+0.10}_{-0.13}$ $1.44^{+0.13}_{-0.10}$ $2.47^{+0.19}_{-0.14}$ $6.08^{+0.30}_{-0.30}$ Ponomareva et al. (2023)
AUDS 0.16 $10.14\pm 0.09$ $1.37\pm 0.05$ $2.55\pm 0.01$ $3.55\pm 0.29$ Hoppmann et al. (2015)
BUDHIES 0.2 $9.77^{+0.16}_{-0.04}$ $1.49\pm 0.48$ $2.25\pm 0.03$ $4.10\pm 4.43$ Gogate (2022)
HI int. mapping 0.32 $9.46^{+0.38}_{-0.31}$ $1.29^{+0.36}_{-0.39}$ $1.80^{+0.65}_{-0.72}$ $4.13^{+5.55}_{-1.58}$ Paul et al. (2023)
Combined (poly.) 0.37 $10.16^{+0.08}_{-0.08}$ $1.48^{+0.10}_{-0.09}$ $2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$ This work
Combined (linear) 0.37 $10.13^{+0.09}_{-0.09}$ $1.56^{+0.14}_{-0.13}$ $2.31^{+0.15}_{-0.17}$ $9.98^{+1.78}_{-1.12}$ This work
Table 3: Numerical results for the best-fitting parameters to the HIMF using a Schechter model as in Eq. 3, for the results from this work and from different literature results. The first column reports the surveys the results refer to; the second column reports the median redshift at which the HIMF has been measured; the third, fourth and fifth columns report the best-fitting values for the $\log_{10}(\tilde{M}/{\rm M}_{\odot})$, $\alpha$, and $\log_{10}(\tilde{\Phi}/{\rm Mpc}^{-3})$; the sixth column reports the inferred value for $\Omega_{\rm HI}$; the seventh column displays the reference presenting the results.
@@
Fig. 2 shows the resulting HIMF from this work as yellow stars (linear model) and blue squares (polynomial model). We display as blue and yellow solid lines the best-fitting Schechter profiles as found from the median of the posterior distributions of the parameters. We also display the results for the HIMF from the literature reported in Table 3. In particular, we show the results from ALFALFA at $z\sim 0$ (Jones et al., 2018, light green dashed), from MIGHTEE at $0<z<0.084$ (Ponomareva et al., 2023, red dotted and gray dotted-dashed), from the AUDS (Hoppmann et al., 2015, brown dotted-dashed) and BUDHIES (Gogate, 2022, cyan dashed) surveys at $z\sim 0.16$ and $z\sim 0.2$, respectively (all based on direct detections), from a HI intensity mapping experiment at $z\sim 0.32$ (Paul et al., 2023, purple dotted-dashed), and green circles and orange the semi-empirical HIMF results obtained by Bera et al. (2022) at $z\sim 0.35$ and Chowdhury et al. (2024) at $z\sim 1$ following similar approaches to the one adopted herein. We find difference between the HIMF reported in this work and the one at $z\sim 0$ from ALFALFA. In particular, the HIMF derived here with both models have a consistently higher normalization than the ALFALFA HIMF, on average by a factor $\sim 2$. Compared to the MIGHTEE results at lower redshift and to the AUDS survey at $z\sim 0.16$, the HIMF from this work has a larger normalization at low $M_{\rm HI}$ values, whereas they tend to converge towards high $M_{\rm HI}$. The results from BUDHIES, from Bera et al. (2022), and from Paul et al. (2023) have comparable normalization to the one of our results at low $M_{\rm HI}$, but they present a rapid fall-off towards high $M_{\rm HI}$, dropping below the ALFALFA HIMF at $\log_{10}(M_{\star}/{\rm M_{\odot}})\lesssim 9.6$. Essentially, there is evidence for a variation between our results at $z\sim 0.37$ and the results by Chowdhury et al. (2024) at $z\sim 1$, the latter having a consistently larger normalization at the probed H I masses. These differences could be due to potential evolution of the HIMF over cosmic times, or to possible selection biases or other systematics. In particular, the approach used in this work relies on a galaxy sample which is extracted from the COSMOS2020 catalog, and hence, selected in optical $M_{\star}$. This is not the case for e.g. blind H I surveys to which we are comparing our results. In general, the mentioned H I surveys feature a severe inhomogeneity in the covered area, depth and sensitivity, instrument and observational configuration (e.g., single dish vs interferometric), among others. In addition, potential systematics related to the interferometric data from MIGHTEE and CHILES may also have an impact in the final results. Finally, as mentioned earlier in this section, the results on the HIMF to which we are comparing our findings are obtained following different methods. Therefore, given these aspects altogether, it is hard to conclusively tell if the detected differences between our results and the other tabulated results are fully due to cosmic evolution or are contaminated by the effects discussed above.
@@
  • the resulting values for $\Omega_{\rm HI}$ obtained by assuming the linear and the polynomial model for the scaling relation differ at $\sim 2.4\sigma$ for the combined stacking;
@@
Figure 2: Compilation of HIMF results from this work and from the literature. We show as yellow stars and blue squares the resulting HIMF from the linear and polynomial models, respectively, as well as their best-fitting models as solid lines of the same color. We display the HIMF from ALFALFA at $z\sim 0$ (Jones et al., 2018) as a green dashed line, from MIGHTEE at $0<z<0.084$ as red dotted ($1/V_{\rm max}$) and gray dotted-dashed (MML) lines (Ponomareva et al., 2023), from AUDS at $z\sim 0.16$ as a brown dotted-dashed line (Hoppmann et al., 2015), from BUDHIES at $z\sim 0.2$ as a cyan dashed line (Gogate, 2022), from the intensity mapping experiment by at $z\sim 0.32$ as a purple dotted-dashed line (Paul et al., 2023), and as green circles and orange diamonds the results from a similar approach as the one followed herein applied to stacking results from uGMRT data at $z\sim 0.35$ (Bera et al., 2022) and $z\sim 1$ (Chowdhury et al., 2024).